OpenStudy (anonymous):

Freddie is at chess practice waiting on his opponent's next move. He notices that the 4-inch-long minute hand is rotating around the clock and marking off time like degrees on a unit circle. Part 1: How many radians does the minute hand move from 3:35 to 3:55? (Hint: Find the number of degrees per minute first.) Part 2: How far does the tip of the minute hand travel during that time? Part 3: How many radians on the unit circle would the minute hand travel from 0 degrees if it were to move 3pi inches? Part 4: What is the coordinate point associated with this radian measure?

OpenStudy (anonymous):

I got the first part I just need help on the second. The answer for the first part is 2π/3 radians

OpenStudy (anonymous):

@jim_thompson5910 ?

jimthompson5910 (jim_thompson5910):

I'm getting 2pi/3 as well

jimthompson5910 (jim_thompson5910):

what is the circumference of the entire circle?

OpenStudy (anonymous):

360°

jimthompson5910 (jim_thompson5910):

that's the number of degrees in a full rotation

jimthompson5910 (jim_thompson5910):

I want the perimeter of the circle

OpenStudy (anonymous):

It doesn't give one. I can only think of 360

jimthompson5910 (jim_thompson5910):

formula C = 2*pi*r

OpenStudy (anonymous):

This is the circle

jimthompson5910 (jim_thompson5910):

the clock's minute hand has a length of 4 so r = 4

jimthompson5910 (jim_thompson5910):

I'm talking about the minute hand's circle, not the unit circle

OpenStudy (anonymous):

I got 25.12

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's the distance around the whole circle

jimthompson5910 (jim_thompson5910):

but we only want a piece of that circle (from 35 min to 55 min)

OpenStudy (anonymous):

Okay how would we find that?

jimthompson5910 (jim_thompson5910):

how did you find 2pi/3 from the previous part?

OpenStudy (anonymous):

\[360\div60=6\] \[6\times20=120\] so 120° on the circle has 2π/3 radians

jimthompson5910 (jim_thompson5910):

good, so 120 degrees is 120/360 = 1/3 of the full circle

jimthompson5910 (jim_thompson5910):

multiply 1/3 by the circumference to get the distance just from 35 to 55

OpenStudy (anonymous):

I got 8.37 (rounded)

jimthompson5910 (jim_thompson5910):

me too

jimthompson5910 (jim_thompson5910):

that's the approximate distance around the circle edge from 35 to 55

OpenStudy (anonymous):

Okay thank you :) could you help me with part three too?

jimthompson5910 (jim_thompson5910):

sure

jimthompson5910 (jim_thompson5910):

The exact circumference is C = 2*pi*4 = 8pi agreed?

OpenStudy (anonymous):

Yes

jimthompson5910 (jim_thompson5910):

8pi inches is a full circle the minute hand travels 3pi inches what fraction of the circle does it travel?

OpenStudy (anonymous):

8pi/3pi?

jimthompson5910 (jim_thompson5910):

it's the other way around 3pi/8pi = 3/8

jimthompson5910 (jim_thompson5910):

so it travels 3/8 around the circle

jimthompson5910 (jim_thompson5910):

3/8 of 2pi radians = ???

OpenStudy (anonymous):

\[\frac{ 2π }{ 1 }\times \frac{ 3 }{ 8 }?\] \[\frac{ 6π }{ 8 }=\frac{ 3π }{ 4 }?\]

jimthompson5910 (jim_thompson5910):

that is the radian measure

OpenStudy (anonymous):

Oh and the points would be\[(\frac{ -√2 }{ 2 }, \frac{ √2 }{ 2 })\]

jimthompson5910 (jim_thompson5910):

very close

jimthompson5910 (jim_thompson5910):

keep in mind that the radius is not 1 it's 4

OpenStudy (anonymous):

I thought we were looking for the radian measure. The next part is the points of that radian measure.

jimthompson5910 (jim_thompson5910):

yes it's points on a circle with radius 4